OpenStudy (anonymous):

Is the following set of forces acting on an object able to produce equilibrium ? squareroot5N, 6N, 9N

OpenStudy (aravindg):

is that a math question?

OpenStudy (anonymous):

yup, grade 12 calculus

OpenStudy (anonymous):

i was given if a+b+c=0 then a,b,c are in a state of equilibrium if that helps, i dont know how to apply it.

OpenStudy (klimenkov):

\(\sqrt5N,6N,9N\) What is N?

OpenStudy (anonymous):

i dont think it holds a value, i think it's just there to tell me that the numbers are forces

OpenStudy (amistre64):

i would say those forces depend on how they are being applied

OpenStudy (amistre64):

N is a Newton of Force

OpenStudy (klimenkov):

Ok. So the question actually is "Is there a triangle with sides \(\sqrt5,6,9\)?"

OpenStudy (anonymous):

i think i am supposed to put them in a triangle to figure out if they could produce equilibrium

OpenStudy (amistre64):

a triangle and equilibrium are different concepts

OpenStudy (amistre64):

if anything, all those forces are pulling on a single object, and is there anyway that the directions that they are pulling create a state where the object itself is not moving

OpenStudy (anonymous):

but don't i put the forces together in a triangle to find the resultant, then the opposite of the resultant is the equilibrium ?

OpenStudy (amistre64):



OpenStudy (amistre64):

if we play tug of war with the 6 and 9, the 9 is winning by 3

OpenStudy (anonymous):

so it could produce equilibrium ?

OpenStudy (amistre64):

is there a way to configure the arrows of force such that all directions cancel out?

OpenStudy (anonymous):

i think that's what im supposed to do, and if it cancels out i know it can produce equilibrium but i dont know how

OpenStudy (amistre64):



OpenStudy (amistre64):

a+b = 9 and the up and down force would have to cancel out if we could determine a suitable setuop

OpenStudy (amistre64):

sin(\(\alpha\))=c/6 sin(\(\beta\))=c/sqrt(5) 6sin(\(\alpha\))=c=\(\sqrt{5}\)sin(\(\beta\)) 6sin(\(\alpha\))-\(\sqrt{5}\)sin(\(\beta\))=0 6cos(\(\alpha\))+\(\sqrt{5}\)cos(\(\beta\))=9 hmm

OpenStudy (anonymous):

are you using sin law? sorry im not following

OpenStudy (amistre64):

im using some trig notions yes, but we might be able to do this without all that. first, is 6+sqrt(5) >= 9 ?

OpenStudy (anonymous):

nope, 8.2

OpenStudy (amistre64):

then the answer is no. if both 6 and sqrt(5) are pulling with all their might, 9 still wins if 6 and sqrt(5) pull at different angles, their combined efforts in the direction against 9 is even smaller still.

OpenStudy (anonymous):

oh so if a+b = c it is in a state of equilibrium ?